Calculate the anion gap.

Na 145 mEq/L

K 3.9 mEq/L

Cl 100 mEq/L

HCO3 20 mEq/L

BUN 28 mg/dL

Cre 1.2 mg/dL

Glu 105
Answer #1

Answer -

Na = 135 - 145 meq /liter

K = 3.5 - 5.1 meq/liter

CL = 98 - 107 meq/liter

HCO3 = 22 - 29 meq/liter

BUN= 8 - 25 mg /dl

Cre = 0.6 - 1.3 mg/dl

Glu = 70 - 110 mg /dl

Thank you